n(-N: (1/1)*2+(1/2)*3+...+(1/n)(n+1)=n/(n+1)
1
Expert's answer
2012-10-18T07:59:03-0400
1/(k*(k+1))=1/k-1/(k+1)
so
1/(1*2)+1/(2*3)+....+1/(n*(n+1))=1-1/2+1/2-1/3+...+1/(n-1)-1/n+1/n-1/(n+1)=1-1/(n+1)=n/(n+1)